### v2

# OnPrem.LLM: A Privacy-Conscious Document Intelligence Toolkit

Abstract
We present OnPrem.LLM, a Python-based toolkit for applying large language models (LLMs) to sensitive, non-public data in offline or restricted environments. The system is designed for privacy-preserving use cases and provides prebuilt pipelines for document processing and storage, retrieval-augmented generation (RAG), information extraction, summarization, classification, and prompt/output processing with minimal configuration. OnPrem.LLM supports multiple LLM backends—including llama.cpp, Ollama, vLLM, and Hugging Face Transformers—with quantized model support, GPU acceleration, and seamless backend switching. Although designed for fully local execution, OnPrem.LLM also supports integration with a wide range of cloud LLM providers when permitted, enabling hybrid deployments that balance performance with data control. A no-code web interface extends accessibility to non-technical users.
Keywords: generative AI, large language models, LLM, NLP, machine learning
1 Introduction
Large Language Models (LLMs) such as GPT-4, LLaMA, Mistral, and Claude have significantly advanced the state of natural language processing, achieving strong performance across tasks including text generation, summarization, question answering, and code synthesis (Anthropic (2024); Jiang et al. (2023); OpenAI et al. (2024); Hugo Touvron et al. (2023)). While many of these models are accessible through cloud-based APIs, their application to sensitive, non-public data remains challenging in regulated or restricted environments.
Organizations in domains such as defense, healthcare, finance, and law often operate under strict data privacy and compliance requirements. These constraints frequently prohibit the use of general-purpose external services, particularly in environments with firewalls, air-gapped networks, or classified (or otherwise sensitive) workloads.
Common approaches to LLM deployment under these constraints include: (1) running lightweight, local models with reduced performance; (2) self-hosting larger models at significant infrastructure cost; or (3) implementing complex, hybrid pipelines to leverage cloud APIs without violating data governance. Each path incurs trade-offs in accuracy, scalability, latency, and maintainability. Prior work on privacy-focused systems (e.g., PrivateGPT, LocalGPT, GPT4All) typically lack comprehensive end-to-end pipelines for common tasks, do not support a diverse range of LLM backends, offer minimal no-code options for non-technical users, and are often constrained by a narrow focus on specific retrieval strategies (Anand et al. (2023); Iván Martínez (2023); PromptEngineer (2023)).
To address this gap, we introduce OnPrem.LLM, a modular, production-ready toolkit for applying LLMs to private document workloads in constrained environments. The system supports both fully local execution and secure integration with privacy-compliant cloud endpoints, giving organizations flexible control over data locality and model placement. It includes a suite of prebuilt pipelines for common document intelligence tasks such as advanced document processing (e.g., table extraction, OCR, markdown conversion), retrieval-augmented generation (RAG), information extraction, text classification, semantic search, and summarization.
OnPrem.LLM supports a range of LLM backends—including llama.cpp for efficient quantized inference, Hugging Face Transformers for broad model compatibility, Ollama for simplified local model orchestration, and vLLM for high-throughput, GPU-accelerated inference. In addition, the system provides optional connectors to privacy-compliant cloud providers (e.g., AWS GovCloud, Azure Government). [^1] OnPrem.LLM was originally named to reflect its exclusive focus on private, local LLMs. Support for cloud-based models was added later to enable hybrid local/cloud deployments and to support privacy-compliant cloud endpoints. While the toolkit prioritizes privacy, publicly accessible cloud LLMs (e.g., OpenAI, Anthropic) can also easily be used for applications involving public or non-sensitive data (e.g., government policy documents, scientific publications), enabling hybrid deployments that balance performance with data control. A unified API enables seamless backend switching, while a no-code web interface allows non-technical users to perform complex document analysis without programming. Key design principles include:
  • $\bullet$
    Data control – Local processing by default; external access is opt-in and configurable
  • $\bullet$
    Deployment flexibility – Operates on consumer-grade machines or GPU-enabled infrastructure
  • $\bullet$
    Ease of integration – Python API, point-and-click web interface, and prebuilt workflows streamline setup and execution
  • $\bullet$
    Real-world focus – Built-in pipelines solve practical document-centric tasks out of the box
OnPrem.LLM is open-source, free to use under a permissive Apache license, and available on GitHub at: https://github.com/amaiya/onprem. The toolkit has been applied to a wide range of use cases in the public sector including horizon scanning of scientific and engineering research, analyses of government policy, qualitative survey analyses, and resume parsing for talent acquisition. Next, we discuss the core modules in the package.
2 Core Modules
OnPrem.LLM is organized into four primary modules that together provide a comprehensive framework for document intelligence:
2.1 LLM Module
The core engine for interfacing with large language models. It provides a unified API for working with various LLM backends including llama.cpp, Hugging Face Transformers, Ollama, vLLM, and a wide-range of cloud providers (Anthropic (2024); Georgi Gerganov et al. (2023); Ollama Contributors (2025); Kwon et al. (2023); OpenAI et al. (2024); Wolf et al. (2020)). This module abstracts the complexity of different model implementations through a consistent interface while handling critical operations such as model loading with inflight quantization support, easy accessibility to LLMs served through APIs, agentic-like RAG, and structured LLM outputs.
2.2 Ingest Module
A comprehensive document processing pipeline that transforms raw documents into retrievable knowledge. It supports multiple document formats with specialized loaders, automated OCR for image-based text, and extraction of tables from PDFs. The module offers three distinct vector storage approaches:
  • 1.
    Dense Store: Implements semantic search using sentence transformer embeddings and ChromaDB for similarity-based retrieval using HNSW indexes (Huber et al. (2025)).
  • 2.
    Sparse Store: Provides both on-the-fly semantic search and traditional keyword search through Whoosh [^2] We use whoosh-reloaded (available at https://github.com/Sygil-Dev/whoosh-reloaded), a fork and continuation of the original Whoosh project. with stemming analyzers and a rich field schema
  • 3.
    Dual Store: Combines both approaches by maintaining parallel stores, enabling hybrid retrieval that leverages both semantic similarity and keyword (or field) matching
2.3 Pipelines Module
Pre-built workflows for common document intelligence tasks with specialized submodules:
  • $\bullet$
    Extractor: Applies prompts to document units (sentences/paragraphs/passages) to extract structured information with Pydantic model validation (Colvin et al. (2025))
  • $\bullet$
    Summarizer: Provides document summarization with multiple strategies including map-reduce for large documents and concept-focused summarization [^3] Concept-focused summarization is a technique available in OnPrem.LLM to summarize documents with respect to a user-specified concept of interest.
  • $\bullet$
    Classifier: Implements text classification through scikit-learn wrappers (SKClassifier), Hugging Face transformers (HFClassifier), and few-shot learning with limited examples (FewShotClassifier) (Pedregosa et al. (2011); Tunstall et al. (2022); Wolf et al. (2020))
  • $\bullet$
    Guider: Constrains LLM outputs to enforce structural consistency in generated text
2.4 App Module
As shown in Figure 1, a Streamlit-based web application makes the system accessible to non-technical users through five specialized interfaces (Streamlit Contributors (2025)). The web interfaces offer easy, point-and-click access to: 1) interactive chat with conversation history, 2) document-based question answering with source attribution to mitigate hallucinations, 3) custom prompt application to individual document passages with Excel export capabilities, 4) keyword and semantic search with filtering, pagination, and result highlighting, and 5) an administrative interface for document ingestion, folder management, and application configuration. For more information, refer to the web UI documentation. [^4] https://amaiya.github.io/onprem/webapp.html
3 Usage Examples
The package documentation provides numerous practical examples and applications. [^5] https://amaiya.github.io/onprem/ To illustrate the software’s ease of use, Example 1 presents a concise example of retrieval-augmented generation (RAG). We download and ingest the House Report on the 2024 National Defense Authorization Act (NDAA) to answer a related policy question.
Example 1: A Basic RAG Pipeline in OnPrem.LLM
from onprem import LLM, utils
# STEP 1: download the 2024 NDAA report
url = 'https://www.congress.gov/118/crpt/hrpt125/CRPT-118hrpt125.pdf'
utils.download(url,'/tmp/ndaa2024/report.pdf')
# STEP 2: load the default LLM using the default LLM backend
llm = LLM(n_gpu_layers=-1)
# STEP 3: ingest documents into default vector store
llm.ingest('/tmp/ndaa2024')
# STEP 4: ask questions
result = llm.ask('What is said about hypersonics?')
[…]
ANSWER:
The context provided highlights the importance of expanding programs related to
hypersonic technology. The House Committee on Armed Services has directed the
Secretary of Defense to (answer truncated due to space constraints)
4 Conclusion
OnPrem.LLM addresses the critical need for privacy-preserving document intelligence in restricted environments. By combining local large language model (LLM) inference, along with cloud-based LLM options, and a modular architecture with prebuilt pipelines, the toolkit enables organizations to implement advanced NLP workflows without compromising data governance. Its design allows for flexible control over the system footprint, making it suitable for a variety of application environments. As LLMs continue to improve in capability and efficiency, the demand for frameworks that support privacy-conscious, resource-aware use will only grow.
Figure 1: The OnPrem.LLM Web UI.